OpenStudy (anonymous):

Solve using elimination: 6x − 2y − 3= 0 4y − 4x = -6 Select one: a. (0, -3/2) b. (2, 9/2) c. (7, 4) d. (-8,-5)

OpenStudy (gorv):

4y-4x=-6 add 4x on both side and tell what u got

OpenStudy (kropot72):

First arrange both equations in standard form, giving us: 6x - 2y = 3 ...........(1) -4x +4y = -6 ........(2) If we multiply both sides of equation (1) by 2, we will make the coefficient of y equal and opposite to the coefficient of y in equation (2). Therefore we will have these two equations: 12x -4y = 6 .........(3) -4x +4y = -6 ........(2) If equations (3) and (2) are now added together, we will eliminate the terms in y and be able to find the value of x. Can you do that?

OpenStudy (kropot72):

@gorv Sorry, I did not know you were helping with this.

OpenStudy (anonymous):

14x?

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

12x -4y = 6 .........(3) -4x +4y = -6 ........(2) If we add equations (3) and (2) we get: 12x - 4x -4y + 4y = 6 - 6 ..............(4) Now you need to simplify equation (4) by combining like terms. Can you do that?

OpenStudy (anonymous):

so times 4 by 12?

OpenStudy (anonymous):

8x=0 ?

OpenStudy (kropot72):

Yes, you are correct. 8x = 0, therefore x = 0. Now if you plug the value of x into equation (2) you can find the value of y.

OpenStudy (anonymous):

so then a? @kropot72

OpenStudy (kropot72):

Yes, a is the correct choice.

OpenStudy (anonymous):

gracias @kropot72

OpenStudy (kropot72):

You're welcome :)